# Self-Rewarding Vision-Language Model via Reasoning Decomposition

### v1

@@ 3.3.4 Analysis on Text-only Reasoning
4 Related Work
4.1 Post-Training Vision-Language Models
Recent vision-language models have increasingly leveraged post-training alignment techniques, including instruction tuning and reinforcement learning, to enhance general-purpose multimodal performance (Liu et al., 2023b; Bai et al., 2025; Chen et al., 2024; et al, 2024; Huang et al., 2025). For example, LLaVA (Liu et al., 2023b) is tuned on GPT-4 generated (image, question, answer) pairs, coupling a CLIP encoder with Vicuna to produce a visual chat assistant that imitates some GPT-4 vision capabilities. InstructBLIP (Dai et al., 2023) introduces an instruction-aware query transformer tuned on 26 datasets, which yields a model that substantially outperforms even larger models on zero-shot benchmarks. Beyond standard instruction-tuning methods like LLaVA and InstructBLIP, recent work increasingly uses reinforcement learning (RL) to align vision-language models for better reasoning (Huang et al., 2025; Xia et al., 2025; Xiao et al., 2025). Many of these methods, inspired by techniques from DeepSeek-R1 (DeepSeek-AI et al., 2025), focus on sophisticated reward engineering. Strategies include providing step-wise rewards to supervise the intermediate reasoning (Zhang et al., 2025), adding explicit visual annotations to ground truth for calculating visual rewards (Xiao et al., 2025), and applying RL in a two-stage curriculum that first strengthens text-only reasoning (Peng et al., 2025b). As a complementary approach, RL from AI Feedback for VLMs demonstrates that preference-based alignment is also a powerful signal, showing it can substantially reduce object hallucination by learning from AI-generated feedback (Yu et al., 2024a).
4.2 Self-Rewarding Reinforcement Learning
The existing reinforcement learning with verifiable rewards (RLVR) methods heavily rely on high-quality reward models or human feedback, creating a major bottleneck for scalability (Peng et al., 2025a; Dai et al., 2025; Li et al., 2025c; Luu et al., 2025). To overcome this, recent work explores self-rewarding approaches, where the model itself provides intrinsic reward signals during RL post-training, an idea first pioneered by Yuan et al. (2025). Building on self-rewarding language models, methods replace external reward models with the model’s own confidence and uncertainty (logit-based self-certainty) or self-verification of its solutions, and even elicit a latent endogenous reward already present inside base LLMs (Zhao et al., 2025; Li et al., 2025a; Simonds et al., 2025; Zheng et al., 2025; van Niekerk et al., 2025). For example, RLIF leverages self-certainty as a reward, achieving comparable performance to GRPO while improving out-of-distribution generalization (Zhao et al., 2025). Similarly, RLSC optimizes a self-confidence reward to secure large accuracy gains with only a few training samples (Li et al., 2025a).
Although self-generated reward signals have thrived in text-only LLMs, only a few works extend this idea to VLMs (Zhou et al., 2024; Lee et al., 2025; Holmes & Chi, 2025), largely due to the complexity of the visual modality and the difficulty of properly defining and evaluating reward signals that capture visual perception. Recent progress includes Calibrated Self-Rewarding, which iteratively generates candidates, self-scores them with step-wise, visually constrained rewards, and fine-tunes via direct preference optimization (DPO) (Zhou et al., 2024). Similarly, RG-VLM uses a VLM to directly label rewards for offline trajectories in long-horizon visual tasks, serving as an auxiliary signal that boosts generalization (Lee et al., 2025). Beyond judgment-based signals, ARES derives dense shaped rewards from attention weights to accelerate learning under sparse or delayed feedback (Holmes & Chi, 2025). These works show that internal visual signals can provide rich reward feedback for VLM alignment without costly supervision, yet the reward is not integrated end-to-end, where the policy receives both visual perception and answer rewards during training.
5 Conclusion and Future Work
In this paper, we introduce Vision-SR1, a self-rewarded reinforcement learning framework that decomposes vision-language understanding into visual perception and language reasoning by explicitly rewarding visual perception by VLM itself. Vision-SR1 strengthens visual perception and reduces language shortcuts, thereby improving VLM performance across several domains of vision-language tasks. Our proposed metric LSR further shows how perception reward lowers the tendency of models to answer via language shortcut rather than genuine visual reasoning.
This work opens up several future research directions:
First, future research can explore more explicit perception rewards such a rewarding visual embeddings in answering questions directly, rather than converting them into textual cues to reduce information loss and lead to more robust and grounded visual understanding.
Second, exploring self-evolving VLMs that reward both visual perception and answers without using external signals (e.g. human labels) that can enhance VLM’s general visual reasoning abilities.
It is also important to recognize that some of the observed mathematical gains from RL training in VLMs may come from spurious effects – for instance, recalibrating the LLM backbone’s output distribution can boost multimodal math performance without true visual grounding. This suggests that improvements in accuracy may sometimes reflect better exploitation of language shortcuts rather than genuine perception gains. Therefore, future work can also explore more analysis to disentangle visual grounding from shortcut learning. Establishing a stronger foundation for measuring and mitigating language shortcuts (benchmarks) will enable the community to design more principled training objectives and ultimately build VLMs with deeper, more reliable visual reasoning.
Acknowledgements
We thank Hongming Zhang, Linfeng Song, Ruosen Li, Zhiyuan Ren for their helpful discussions!
Appendix A Experiment Details
A.1 Prompt Templates
[…]
A.2 LLM-as-a-Judge Prompt
We use Gemini-2.5-flash as our LLM-as-a-Judge to evaluate
Question: {Question}

### v2

@@ B.2.1 Analysis on Text-only Reasoning
B.2.2 Ablation Study on Self-Reward
We train a control version of our model without the visual reasoning self-reward and Multi-Reward Policy Optimization (Vision-SR1 w/o self-reward). This ablated model still follows a structured output (visual perception, CoT reasoning, and answer) but is optimized only with answer and format rewards. The self-visual reward for self-evaluating visual reasoning and Multi-Reward Policy Optimization are removed. We note the only difference between Vision-SR1 w/o self-reward and Vision-R1 (Huang et al., 2025b) lies in the output structure, i.e., using different system prompts, while all supervision signals (answer reward and format rewards) remain the same. Interestingly, our system prompt yields slightly better performance (+1.0 on average). Table 6 reports the ablation results. We find that not including visual reasoning reward and Multi-Reward Policy Optimization could lead to overall worse VLM task performance compared to including them in the training process.
Table 6: Results of ablation study: Vision-SR1 v.s. Vision-SR1 w/o visual perception self-reward.
General Visual Understanding Visual Math & Hallucination
Methods MMMU MMMU RealWorld VisNum Math MATH Hallusion Avg.
-Pro QA Bench Verse -Vision Bench
Vision-SR1 (3B) 40.8 49.6 66.1 41.9 45.8 29.3 68.3 48.8
  $\vdash$ w/o self-reward 40.0 48.0 62.6 41.6 45.1 30.2 65.8 47.6
Vision-SR1 (7B) 40.7 52.2 69.2 43.5 54.5 36.2 68.9 52.2
  $\vdash$ w/o self-reward 42.8 51.8 67.3 35.7 52.6 34.4 67.8 50.3